Solve by completing the square:
x^2 + 6x + 1 = 0

Answer by nerdybill(7384)

This site describes it quite well:
http://www.purplemath.com/modules/solvquad3.htm
.
x^2 + 6x + 1 = 0
.
First, isolate the x's -- so, subtract 1 from both sides:
x^2 + 6x = -1
.
Next, look at the coefficient associated with the x term (6). Divide this by 2 and square it to get 9.
.
So, we ADD 9 to both sides:
x^2 + 6x + 9 = -1 + 9
.
Now, we factor the left side and combine the right to get:
(x+3)(x+3) = 8
(x+3)^2 = 8
.
To solve:
(x+3)^2 = 8
Take the square root of both sides:
x + 3 = (+-)sqrt(8)
x = (+-)sqrt(8) - 3
x = (+-)2sqrt(2) - 3
.
Note: the (+-) means "plus or minus"